### v1

# ICAT: Evaluating Completeness of Factual Information in Long-form Text Generation

Abstract
This paper presents ICAT, an evaluation framework for measuring coverage of diverse factual information in long-form text generation. ICAT breaks down a long output text into a list of atomic claims and not only verifies each claim through retrieval from a (reliable) knowledge source, but also computes the alignment between the atomic factual claims and various aspects expected to be presented in the output. We study three implementations of the ICAT framework, each with a different assumption on the availability of aspects and alignment method. By adopting data from the diversification task in the TREC Web Track and the ClueWeb corpus, we evaluate the ICAT framework. We demonstrate strong correlation with human judgments and provide comprehensive evaluation across multiple state-of-the-art LLMs. Our framework further offers interpretable and fine-grained analysis of diversity and coverage. Its modular design allows for easy adaptation to different domains and datasets, making it a valuable tool for evaluating the qualitative aspects of long-form responses produced by LLMs.
Beyond Factual Accuracy: Evaluating Coverage of Diverse Factual Information in Long-form Text Generation
Chris Samarinas1thanks: Equal contribution.    Alexander Krubner211footnotemark: 1thanks: Work done while visiting the Center for Intelligent Information Retrieval at UMass Amherst.    Alireza Salemi1 Youngwoo Kim1    Hamed Zamani1 1University of Massachusetts Amherst  2Salzburg University of Applied Sciences 1{csamarinas, asalemi, youngwookim, zamani}@cs.umass.edu 2akrubner.bin-m2022@fh-salzburg.ac.at
1 Introduction
Recent advancements in large language model (LLM) development have enabled us to go beyond phrase-level and sentence-level generation and build systems capable of generating fluent and statistically plausible long responses. The evaluation of these systems remains a challenging area of research Samarinas et al. (2024). Numerous works have proposed methods that measure lexical overlap with one or more reference outputs, such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005), as well as semantic similarity-based methods, such as BERTScore Zhang et al. (2019) and G-Eval Liu et al. (2023), for long-form text generation. One major issue with all these methods is that many different long outputs by LLMs can be considered equally acceptable; given the size of the output space in long-form text generation, it is impractical to create reference texts that can cover all possible long outputs at scale. This has recently motivated researchers to develop reference-free evaluation methodologies, such as Prism Agrawal et al. (2021); Thompson and Post (2020). However, they are often unable to evaluate hallucinations and biases in LLMs. FActScore Min et al. (2023) and VERISCORE Song et al. (2024) address this issue by looking at fine-grained atomic claims in the generated output that can be backed up by retrieved information from a given (reliable) source. These methods solely evaluate the factual accuracy in the generated output.
We argue that factual accuracy is not sufficient when evaluating LLMs. Imagine a user submits the query “is coffee good for health?” to an LLM. Given the current state of knowledge in medicine and related disciplines, there are some health benefits and also some health risks to consuming coffee. If the LLM generates an output that only covers the health benefits of coffee by making factual and grounded arguments, it would achieve a perfect FActScore and VERISCORE. Therefore, answer completeness and coverage of diverse factual information should be a major component when evaluating LLMs for long-form text generation, such as non-factoid and complex comparative question answering, report generation, etc.
This paper introduces ICAT, [^1] ICAT: Information Coverage & Accuracy in Text. Code is publicly available at https://github.com/algoprog/ICAT a reference-free evaluation framework for long-form text generation that considers both factual accuracy and coverage of diverse factual information. Inspired by FActScore Min et al. (2023) and VERISCORE Song et al. (2024), ICAT breaks down the generated long text into atomic claims. Through retrieval from a (reliable) corpus $C$ or the Web, ICAT verifies each atomic claim to ensure its factuality. To measure completeness and coverage of diverse facts, ICAT requires a set of diverse aspects to compute an alignment between each atomic factual claim in the LLM response and the set of diverse aspects. We study three implementations of the ICAT framework as follows: ICAT-M assumes that a ground-truth set of diverse claims are obtained manually and is available to the evaluation framework. It also assumes that the groundtruth relevance annotation for each document in the corpus $C$ to each aspect is provided. Using this information, the retrieval model can identify which aspect is being covered by each atomic factual claim in the LLM response. ICAT-S similarly assumes that a groundtruth set of diverse claims are obtained manually, however no aspect-level relevance judgment is available. Therefore, it uses an LLM to conduct pseudo-labeling and performing alignment between the atomic factual claims and the set of aspects. On the other hand, ICAT-A assumes that the aspect set is not available, so it first uses an LLM to automatically generate diverse aspects of the input and then conduct pseudo-labeling for alignment, as is done in the second variant.
In our experiments, we use ClueWeb The Lemur Project (2009) as the retrieval corpus. We solely focus on the English documents of the ClueWeb collection. For web-based grounding, we use the Brave Search API. For experiments, we rely on the input topics from the TREC Web Track Clarke et al. (2009, 2010, 2011, 2012). The argument for this decision is based on the fact that TREC Web Track topics have also been used for search result diversification. This means that the topics include up to 7 subtopics (aspects) and documents are provided with aspect-level relevance annotations. Our experiments show that there is strong corelation of ICAT with human judgments (ICAT-S and ICAT-A achieving Pearson’s $\rho>0.4$), showcasing the utility of this framework for evaluating coverage of diverse factual information in LLM responses without human input.
By offering a modular and adaptable framework, ICAT enables researchers to tailor the evaluation process to specific needs, making it a valuable tool for assessing the qualitative aspects of long-form responses produced by LLMs. The decomposition of LLM outputs into atomic claims and their alignment with specific topics makes the evaluation process highly interpretable - evaluators can trace exactly which claims support which topics and identify gaps in coverage. This granular analysis capability, combined with the framework’s ability to evaluate both factual accuracy and topic coverage, provides a more comprehensive assessment compared to existing metrics that only measure one of these aspects. ICAT’s implementation is available at https://github.com/algoprog/ICAT.
Figure 1: Retrieval-based evaluation of LLM responses with ICAT. Topic generation and coverage models are optional depending on the chosen evaluation method.
2 Related Work
Text Generation Evaluation Traditional approaches to evaluating generated text have primarily focused on n-gram overlap metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005). While these metrics are effective for assessing local coherence and fluency, they fail to capture higher-level aspects such as topic coverage and diversity. Recent work has introduced more sophisticated metrics like BERTScore Zhang et al. (2019), BLEURT Sellam et al. (2020), and unified multi-dimensional evaluators Zhong et al. (2022) which leverage pre-trained LLMs for more nuanced evaluation.
Topic Coverage and Diversity Research on evaluating topic coverage has roots in information retrieval, where metrics like $\alpha$-nDCG Clarke et al. (2008) and S-Recall were used to assess the topical diversity of search results. The concept of diversity in evaluating generated text can encompass various interpretations, including lexical diversity (analyzing the variety of words used) and topical diversity (assessing the range of topics covered). In the context of text generation, recent work has explored various approaches to measuring lexical diversity, including term overlap self-similarity such as Self-BLEU Zhu et al. (2018) and the proportion of distinct unigrams and bigrams in generated responses Li et al. (2016). However, research on evaluation of topical diversity in LLMs is currently limited.
LLM Evaluation Frameworks Several frameworks have been proposed for evaluating different aspects of LLM performance, including factuality Min et al. (2023); Song et al. (2024) and dialogue quality Mehri and Eskenazi (2020). Concurrent to this research, the AutoNuggetizer framework Pradeep et al. (2024) used LLMs to generate and assess the coverage of nuggets in text. However, in that framework, the nuggets were specific claims rather than higher-level subtopics. Moreover, AutoNuggetizer solely focuses on nugget coverage and does not consider factuality of responses.
Our work builds upon these foundations, specifically addressing the challenge of evaluating topic coverage in long-form text generation while considering factuality at the same time.
3 ICAT
Queries that require a long-form response, e.g., complex non-factoid questions, are often associated with multiple aspects. The response to these queries often include multiple claims, some of which may be factually accurate, while others may be inaccurate. An ideal response to these queries should not only contain factually accurate claims, but should also leave no aspect or perspective unaddressed. For instance, an ideal answer to a question about a legislation should cover perspectives from all political parties. An ideal answer to a question about the impact of a food or a medication on health should cover both positive, neutral, and negative perspectives. However, no existing evaluation metric can evaluate both factual accuracy and aspect coverage in long-form text generation. To address these, given a long output $y$ produced in response to an input $x$, the ICAT framework computes two main scores: factuality score and coverage score.
Factuality Score.
Building upon prior work, such as FActScore Min et al. (2023) and VERISCORE Song et al. (2024), Factuality Score measures the ratio (or percentage) of factually accurate claims in $y$. To do so, it is crucial that the generated claims are accurate. Let $AC(y)$ be a function that extracts atomic claims from a generated response $y$. Given the set of atomic claims $C=AC(y)$ made in $y$, we define the function $C_{T}=CG(C;K)$ that verifies the factuality of claims in $C$ using a given knowledge source $K$. Therefore, $C_{T}\subseteq C$ denotes the set of factually verified claims in response $y$. Factuality Score is then defined as follows:
$S_{\text{fact}}=\frac{|C_{T}|}{|C|}$
where ||$\cdot$|| ⋅ | denotes the cardinality of the given set.
Coverage Score.
To evaluate information coverage and diversity in $y$, Coverage Score measures the ratio of aspects being covered by the factually accurate claims in the given text. Hence, it is essential to identify which query aspects are accurately addressed in the generated response. Formally, coverage score can be defined as:
$S_{\text{coverage}}=\frac{|\{TO(c,K):c\in C_{T}\}\cap TQ(x)|}{|TQ(x)|}$
where $TO$ is a function that identifies the subtopics associated with claim $c$, and $TQ$ is a function that returns all aspects related to the input $x$. Note that aspect coverage is only computed for factually verified claims, i.e., $C_{T}$, instead of all claims. The reason is that non-factual claims should be avoided, regardless of the aspect they cover. Thus, they should not contribute to the coverage score.
The $\text{ICAT}_{\beta}$ Score.
Inspired by F-measure Van Rijsbergen (1979), we calculate the weighted harmonic average of these factuality and coverage scores, as follows:
$\text{ICAT}_{\beta}=(1+\beta^{2})\frac{S_{\text{fact}}\cdot S_{\text{coverage}}}{\beta^{2}S_{\text{fact}}+S_{\text{coverage}}}$
where parameter $\beta$ is a hyper-parameter that controls the trade-off between the factuality and coverage scores. In more detail, $\beta$ controls the weight of Coverage Score compared to Factuality Score. Thus, a higher $\beta$ signifies the impact of information coverage, while a lower $\beta$ prioritizes factual accuracy. The default value for $\beta$ is equal to $1$, where factuality and coverage score are weighted uniformly. Throughout this paper, when the value of $\beta$ is not explicitly mentioned, the default value of $1$ is being used.
The rest of this section provides details on how to develop models for generating factual claims (i.e., function $AC$), how to validate the factuality of claims (i.e., function $CG$), and how to obtain query aspects and compute an alignment between factual claims and all query aspects (i.e., functions $TQ$ and $TO$). The rest of this section describes the approaches at high level to introduce the generic ICAT, while Section 4 provides the implementation details used in our experiments.
Figure 2: Example of atomic claim generation
3.1 Atomic Claim Generation
The atomic claim generation process seeks to break down a given long text into standalone and atomic claim statements that preserve key context and maintain claim consistency (see Figure 2). The generated claims should strike an appropriate balance of granularity, ensuring they are self-contained and decontextualized. We assume the existence of a function $C=AC(y)$, which returns a set of atomic claims $C$, given the long output text $y$. There are various ways to implement this; one might consider each sentence or paragraph in $y$ as an atomic claim. However, this simple approach does not satisfy our expected self-containment and decontextualization qualities. Instead, we follow Min et al. (2023) and utilize an LLM $M_{\text{claims}}$ with the prompt shown in Figure 5. This prompt instructs the LLM to decompose the generated response $y$ into multiple self-explanatory and decontextualized sentences, each containing a single atomic fact. These sentences then constitute the set of atomic claims for the generated output, denoted as $C=AC(y)$. An example of this process is illustrated in Figure 2. An instruction-tuned LLM can be used as $M_{\text{claims}}$ in a zero- of few-shot setting; however, we found that smaller-scale LLMs (such as LLaMA Grattafiori et al. (2024) with 8 billion parameters) cannot accurately perform this task without fine-tuning. Therefore, we either use an LLM with higher capacity or distill knowledge into a smaller-scale LLM, enabling faster inference for our framework. The details of this distillation process are provided in section 4.1.
3.2 Claim Grounding
To design the claim grounding function $CG$, for each claim $c\in C$, we employ a retrieval model $R$ to retrieve $n$ documents from the given knowledge source $K$. Subsequently, a natural language inference (NLI) model $M_{\text{NLI}}$ is used to determine whether the claim can be supported by any of the retrieved documents. If the claim can be inferred from at least one of the retrieved documents, it is considered grounded (i.e., validated, thus factually accurate); otherwise, it is not. The function returns a subset of $C$ that are found grounded.
3.3 Aspect Coverage Assessment
To calculate aspect coverage for some given input prompt (query) $x$, it is essential to have a list of diverse aspects for $x$ (i.e., $TQ(x)$) and a method to determine which aspect each claim pertains to (i.e., $TO(c,K):c\in C$).
Methods for Obtaining Diverse Query Aspects ($TQ$):
We propose two main methods to identify all aspects related to the query $x$:
  • Manual–Ground-truth Aspects: In this case, the aspects that should be included in the response to the query $x$ are provided as a reference for evaluation.
  • Automatic–LLM-based Aspect Generation: Building on previous work showing LLMs can effectively identify aspects of a query Samarinas et al. (2022), we use an LLM $M_{\text{subtopic}}$ with the prompt shown in Figure 6. This prompt instructs the LLM to generate up to 10 aspects for the query, covering the key aspects about it. This approach is useful when ground-truth aspects are unavailable.
Methods of Obtaining the Aspects of an Atomic Claim ($TO$):
We use two methods to identify the aspects related to an atomic claim $c$:
  • Manual–retrieval-based method with aspect-level ground-truth alignment: In this method, we assume access to a knowledge source $K$, where each document is annotated with the aspects it covers. To find the aspects that the claim $c$ covers, we use the retrieval model $R$ to retrieve $n$ documents. Then, according to the ranking, we find the first ranked document that supports the claim $c$ using the method in Section 3.2. The aspects of this document are considered as the aspects that the claim $c$ covers. If none of the documents support claim $c$, we assume that it does not cover any query aspect.
  • Automatic–retrieval-based method with aspect-level LLM-based alignment: In this method, we use an aspect-claim alignment LLM $M_{\text{coverage}}$ to determine which aspects each grounded claim covers. Given a query $x$, its aspects $TQ(x)$, and a set of grounded claims $C_{T}$, we prompt the LLM to analyze each claim and identify which aspects it addresses. The prompt (shown in Figure 4) instructs the LLM to output a structured mapping between claims and aspects, where each claim can be mapped to zero, one, or multiple aspects. This approach eliminates the need for aspect-level relevance judgments in the knowledge source while still maintaining a retrieval-based verification of factual accuracy. Unlike the manual method that assumes a claim covers the aspects associated with its supporting document, this method directly analyzes the semantic relationship between claims and aspects, leading to more accurate assessment.
3.4 Variants of ICAT
We study three variants of ICAT implementations based on how they obtain query aspects and compute the alignment between atomic claims and aspects. Table 1 describes the approaches used in these three variants and highlights their differences.
Approach for Obtaining Diverse Aspects Approach for Claim-Aspect Alignment
ICAT-M Manual: ground-truth aspects Manual: retrieval-based method with aspect-level ground-truth alignment
ICAT-S Manual: ground-truth aspects Automatic: retrieval-based method with aspect-level LLM-based alignment
ICAT-A Automatic: LLM-based aspect generation Automatic: retrieval-based method with aspect-level LLM-based alignment
Table 1: The methods used for obtaining query aspects and claim-aspect alignment in each variant of ICAT.
Corpus-based Retrieval Web-based Retrieval
Method Coverage Model Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$ Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$
ICAT-M N/A 0.196 0.184 0.146 0.216 0.204 0.166
ICAT-S Llama-3.1-8B 0.417 0.400 0.314 0.534 0.521 0.442
Llama-3.1-70B 0.489 0.452 0.376 0.612 0.589 0.486
ICAT-A Llama-3.1-8B 0.417 0.400 0.314 0.548 0.529 0.447
Llama-3.1-70B 0.500 0.459 0.382 0.623 0.596 0.492
Table 2: Correlation of information coverage in the proposed evaluation methods and manual human annotations.
4 Implementation Details
4.1 Atomic Claim Generation
The claim generation module was trained with several key objectives in mind, building on recent work in atomic claim extraction Min et al. (2023). The model learned to extract standalone factual statements from text while maintaining factual consistency and simplifying complex statements. Special attention was paid to preserving important context and qualifiers, and generating claims at an appropriate granularity level Song et al. (2024). For this task we used Llama 3.1 8B Grattafiori et al. (2024) fine-tuned using QLoRA Dettmers et al. (2023) on synthetic examples. We tried using models of this size without fine-tuning, however we found that the generated claims are often not de-contextualized properly. Larger models with 70B or more parameters seem to be effective for this task without fine-tuning, however they are very expensive to run, especially for long texts.
The synthetic training data was generated through a multi-stage process by prompting Llama 3.1 405B. We began by generating 200 diverse high-level topics across multiple domains. For each topic, we generated 5 relevant entities. We then created variable-length paragraphs for each entity and generated the associated list of atomic claims for each of them. Using these 1000 synthetic examples, we fine-tuned the model for 1 epoch with batch size 16, learning rate 2e-4 and LoRA parameters $\alpha$ = 16 and rank = 64.
4.2 Topic Generation
Previous works have shown that LLMs can be very effective in query subtopic generation Samarinas et al. (2022). In our framework, for generating ground-truth topics given a query, we use the same base LLM as the one used in claim generation. Here we found that even without fine-tuning, Llama 3.1 8B can produce relevant topics. In order to reduce the need for extra resources to use a base and fine-tuned version of the LLM for claim generation, we use the VLLM library Kwon et al. (2023) to load the base model only once in memory and efficiently serve the adapter for the fine-tuned version.
4.3 Claim Grounding
We implemented a two-stage approach for grounding atomic claims in the given text with a corpus. We first preprocess the corpus and generate chunks for each document with up to 128 words with 32 words overlap. We use a dense embedding model222hf.co/Snowflake/snowflake-arctic-embed-m Merrick et al. (2024) to produce embeddings for all snippets and FAISS Johnson et al. (2019) to build an efficient approximate nearest-neighbor index. We used IVF with HNSW for cluster assignment as our index type for fast search even when providing a large-scale corpus.
In the first stage, a retriever is used to obtain the $k=10$ most relevant snippets in the corpus for each claim. When web search is used instead of a corpus, we use the returned snippets from Brave Search API. In the second stage, a natural language inference model is used to filter only the supported claims. We use a model based on DeBERTa V3 He et al. (2021) fine-tuned on MultiNLI, FEVER and Adversarial NLI Williams et al. (2018); Thorne et al. (2018); Nie et al. (2020). A claim is kept if there is at least one snippet that supports it. Instead of using a LLM for filtering the supported claims, we use a much smaller BERT based model Devlin et al. (2019) fine-tuned specifically for this task.
We limit the snippet length because both NLI and dense embeddings models based on small pre-trained transformer LMs like BERT tend to have lower performance as the input length increases.
4.4 Aspect-Claim Alignment
Topic coverage is assessed using the same base LLM with claim and topic generation. Given a query, a list of enumerated atomic claims and a list of ground truth topics, the LLM is prompted to produce a list of covered topic ids with their associated claim ids in structured jsonl format.
ICAT-M ICAT-S ICAT-A
LLM $S_{\text{fact}}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-M}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-S}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-A}_{1}$
Corpus-based Retrieval
GPT-4 0.343 0.416 0.327 0.453 0.346 0.463 0.354
Llama-3-70B 0.327 0.451 0.335 0.464 0.355 0.466 0.354
Mixtral-8x22B 0.344 0.370 0.297 0.414 0.342 0.409 0.339
Openchat 3.5 (7B) 0.340 0.413 0.329 0.429 0.348 0.424 0.347
Web-based Retrieval
GPT-4 0.363 0.436 0.347 0.473 0.366 0.483 0.374
Llama-3-70B 0.347 0.471 0.355 0.484 0.375 0.486 0.374
Mixtral-8x22B 0.364 0.390 0.317 0.434 0.362 0.429 0.359
Openchat 3.5 (7B) 0.360 0.433 0.349 0.449 0.368 0.444 0.367
Table 3: Evaluation of various LLMs using ICAT. Llama-3.1-70B is used for claim-aspect alignment.
5 Experimental Setup
Dataset.
We conducted our experiments using the ClueWeb09 Category B corpus–a large-scale web collection with over 50 million English documents The Lemur Project (2009). This corpus has been used in TREC Web Track from 2009 to 2012 , consisting of 200 topics, derived from a commercial search engine’s query log, balanced for popularity. Each topic includes a topic title (i.e., often used as the keyword search query), a description (i.e., detailed description of the information need), type, and subtopics (i.e., diverse aspects of the topic). The relevance judgments encompass 38,637 query-document pairs, with 19.06% (7366) marked as relevant. The dataset’s unique advantage lies in its comprehensive coverage of internet content and human-annotated relevance judgments for topical diversity assessment. Relevance was judged either binary or on a five-point scale (later converted to binary), with documents considered relevant when containing useful information for specific subtopics. In our experiments with this collection, we filtered out spam documents using the Waterloo spam scorer Cormack et al. (2011) with the threshold of 70%. We used BM25 to retrieve 1000 documents for each topic (given its title as the query string) and considered these documents for retrieval in our factual verification process. The query set comprises 50 carefully selected faceted queries, each containing 3-8 subtopics representing different aspects of the information need.
Experimental Setup.
We evaluated four state-of-the-art LLMs: GPT-4, Llama-3-70B-Instruct, Mixtral-8x22B-Instruct-v0.1, and Openchat 3.5 (a fine-tune of Mistral-7B) Wang et al. (2023). For each model, we generated responses for each test query. For the baselines in this paper, we used the query descriptions in their original format from the ClueWeb09 dataset as prompts, which are not optimized for producing very diverse outputs.
6 Experimental Results
Human Evaluation Study.
To validate ICAT’s effectiveness, we conducted a comprehensive human evaluation study using Amazon Mechanical Turk (AMT). For each query-answer pair, three independent annotators assessed the coverage of aspects through a custom interface (Figure 7 in Appendix A.4). We limited the HITs to adult workers from the US, UK, Australia and Ireland, with over 98% approval rate who have completed at least 5,000 assignments. The annotators were tasked with identifying whether specific aspects are present in a given LLM-generated text and highlighting corresponding text evidence for each identified aspect. To ensure quality annotations, we provided detailed guidelines with two reference examples. We use majority voting across annotators. The study achieved substantial inter-annotator agreement with Fleiss’s $\kappa$ = 0.829, which is considered as a substantial agreement.
For each query, we calculated Coverage Scores based on the set of covered topics identified by each evaluation method (ICAT variants) and by human annotators, relative to the set of ground truth topics. These per-query coverage scores were then used to compute linear and rank-bsaed correlation metrics (i.e., Pearson’s $\rho$, Spearman’s $\rho$, and Kendall’s $\tau$) between the automated ICAT methods and human judgments.
The correlation analysis between ICAT variants and human judgments (see Table 2) reveals strong performance across most evaluation methods. Using Llama-3.1-70B as the coverage model, ICAT-A achieves the strongest correlations (Pearson’s $\rho$ = 0.500, p < 0.01; Spearman’s $\rho$ = 0.459, p < 0.01). ICAT-S demonstrates similar performance with the same coverage model (Pearson’s $\rho$ = 0.489, p < 0.01; Spearman’s $\rho$ = 0.452, p < 0.01). While computationally efficient, ICAT-M shows weaker correlation (Pearson’s $\rho$ = 0.196, p < 0.05) with human assessments, highlighting the importance of sophisticated coverage models in improving evaluation accuracy.
Comparing factuality and coverage of information in state-of-the-art LLMs using ICAT.
Our experimental results reveal distinct patterns in how different LLMs balance factuality and coverage (see Table 3). Llama-3-70B demonstrates superior Coverage Score (0.451 for corpus-based and 0.471 for web-based retrieval), while GPT-4 and Mixtral-8x22B show comparable factuality score (0.343 and 0.344, respectively). However, Mixtral exhibits notably lower Coverage Score 0.370 compared to GPT-4’s 0.416, resulting in lower overall $\text{ICAT}_{1}$ scores (0.297 vs 0.327). Notably, Openchat 3.5, despite being the smallest model evaluated, maintains competitive performance across all metrics ($S_{\text{fact}}$ = 0.340, $S_{\text{coverage}}$ = 0.413, $\text{ICAT}_{1}$ = 0.329). When employing ICAT-S and ICAT-A with Llama-3.1-70B as the coverage model, we observe improved Coverage Scores across all models. Llama-3-70B maintains its leading position for corpus-based retrieval, with the highest overall $\text{ICAT}_{1}$ score of 0.355, followed closely by GPT-4 (i.e., 0.346). The results suggest that using Llama-3.1-70B as the coverage model produces more conservative but potentially more reliable assessments compared to Llama-3.1-8B, as evidenced by more consistent Coverage Scores across evaluated LLMs.
Controlling the trade-off between factuality and coverage using $\beta$.
Figure 3 illustrates how different values of $\beta$ affect the trade-off between factuality and coverage scores. Users can adjust $\beta$ based on their evaluation priorities: values of $\beta<1$ give more weight to factuality score, while $\beta>1$ emphasizes on coverage. In our experiments, for lower $\beta$ values, GPT-4 and Mixtral demonstrate superior performance in terms of $\text{ICAT}_{\beta}$. However, as $\beta$ increases and coverage becomes more important, GPT-4 and Llama achieve higher scores due to their stronger Coverage Score.
Figure 3: ICAT-Aβ for various LLMs using Llama-3.1-70B as coverage model and the corpus as knowledge source. For low $\beta$ values where factuality has higher weight, GPT-4 and Mixtral have better performance while for higher values of $\beta$, GPT-4 and Llama have higher scores due to higher $S_{\text{coverage}}$.
7 Conclusions and Future Work
We presented ICAT, a comprehensive framework for evaluating topic coverage in LLM-generated text. Through extensive experimentation using the ClueWeb09 dataset, we demonstrated the framework’s effectiveness across different evaluation scenarios, with our best method achieving strong correlation with human judgments. The modular architecture of ICAT enables flexible adaptation to various evaluation requirements, from manual to automatic approaches for aspect identifications and alignment. Our results highlighted several key findings: (1) the importance of sophisticated coverage models in improving evaluation accuracy, (2) the viability of automatic evaluation approaches that maintain comparable performance to methods requiring ground truth annotations, and (3) the framework’s ability to provide meaningful assessments across different LLM architectures and scales.
In future work, individual components of our ICAT could be improved in terms of effectiveness and efficiency. Additionally, exploring the relationship between model size, evaluation accuracy, and computational efficiency could provide valuable insights for practical applications. Last but not least, the potential bias introduced by using the same or similar LLM when generating ground-truth aspects should be investigated. By using our metric, other works can explore methods for optimizing LLMs to produce more comprehensive outputs.
8 Limitations
Our evaluation framework, while showing promising results, suffers from several limitations that should be considered. First, our experiments reveal that even large language models with 70B parameters sometimes struggle with accurate aspect-claim alignment. This suggests that the correlation with human judgments could potentially be improved by specifically optimizing LLMs for this task, either through fine-tuning or more sophisticated prompting strategies.
Second, our current implementation uses zero-shot prompting for query aspect generation without systematic evaluation of this component’s effectiveness. Future work should explore methods to optimize and rigorously evaluate the aspect generation process, potentially through human evaluation or comparison with expert-curated aspect sets. This could lead to more reliable and comprehensive aspect coverage assessment.
Third, there is a potential source of bias when using the same or similar LLM architecture both for generating query aspects and for producing responses for evaluation. This circular dependency might lead to artificially inflated performance metrics if the evaluated model shares similar biases or knowledge patterns with the model used for aspect generation. Future research should investigate the extent of this potential bias and explore methods to mitigate it, such as using diverse model architectures or more comprehensive human-curated aspects for evaluation.
9 Acknowledgments
This work was supported in part by the Center for Intelligent Information Retrieval (CIIR), in part by the Office of Naval Research contract number N000142212688, and in part by NSF grants #2143434 and #2106282. We acknowledge the support from the Austrian Marshall Plan Foundations, Stefan Wegenkittl, and Martin Uray who made Alexander Krubner’s visit to the CIIR possible. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsors.
Appendix A Appendix
A.1 Atomic Claim Generation
Figure 2 illustrates our atomic claim generation process with a concrete example, showing how a complex paragraph is broken down into individual atomic claims while preserving key contextual information.
A.2 Model Performance Analysis
Table 4 presents detailed evaluation results using Llama-3.1-8B as the coverage model. The results demonstrate consistent performance patterns across different retrieval methods (corpus-based vs web search) and highlight the trade-offs between factuality ($S_{F}$) and coverage ($S_{C}$) scores for various models.
LLM Retrieval $S_{\text{fact}}$ ICAT-S ICAT-A
[…]
GPT-4 Corpus 0.343 0.563 0.394 0.563 0.394
Web Search 0.378 0.583 0.414 0.583 0.414
Llama-3-70B Corpus 0.327 0.616 0.401 0.616 0.401
Web Search 0.365 0.636 0.421 0.636 0.421
Mixtral-8x22B Corpus 0.344 0.567 0.401 0.567 0.401
Web Search 0.382 0.587 0.421 0.587 0.421
Openchat 3.5 (7B) Corpus 0.340 0.550 0.391 0.550 0.391
Web Search 0.376 0.570 0.411 0.570 0.411
Table 4: Evaluation results for ICAT-S and ICAT-A (Llama-3.1-8B Coverage Model)
Table 5 analyzes the impact of different retrieval models on the correlation between ICAT variants and human judgments. The results show that more sophisticated dense retrieval models (Snowflake-Arctic-Embed-L) consistently outperform traditional BM25 across all ICAT variants, with improvements particularly notable in ICAT 2 and 3.
A.3 Prompting Details
Figures 4, 5, and 6 provide the detailed prompts used for subtopic coverage assessment, atomic claim generation, and topic generation respectively. These carefully designed prompts are crucial for ensuring consistent and high-quality outputs from the language models used in our framework.
A.4 Human Evaluation Interface
Figure 7 shows our human annotation interface used in the evaluation study. The interface allows annotators to highlight supporting evidence for different aspects of the query in the LLM response, enabling fine-grained assessment of topic coverage. The visualization of multiple annotators’ highlights helps demonstrate the level of agreement in human judgments.
Method Retrieval Model Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$
ICAT-M BM25 0.182 0.175 0.138
Snowflake-Arctic-Embed-M 0.196 0.184 0.146
Snowflake-Arctic-Embed-L 0.205 0.192 0.153
ICAT-S BM25 0.472 0.443 0.368
Snowflake-Arctic-Embed-M 0.489 0.452 0.376
Snowflake-Arctic-Embed-L 0.503 0.461 0.384
ICAT-A BM25 0.483 0.449 0.374
Snowflake-Arctic-Embed-M 0.500 0.459 0.382
Snowflake-Arctic-Embed-L 0.512 0.468 0.389
Table 5: Impact of different retrieval models on correlation with Human Judgments using Llama-3.1-70B for topic-claim alignment.
Figure 4: Subtopic coverage prompt
Figure 5: Claim generation prompt
Figure 6: Topic generation prompt
Figure 7: Human annotation interface showing the supporting evidence highlights from 3 annotators for the given query in the shown LLM response.