### v4

# Beyond Factual Accuracy: Evaluating Coverage of Diverse Factual Information in Long-form Text Generation

Abstract
This paper presents ICAT,111source code: https://github.com/algoprog/ICAT an evaluation framework for measuring coverage of diverse factual information in long-form text generation. ICAT breaks down a long output text into a list of atomic claims and not only verifies each claim through retrieval from a (reliable) knowledge source, but also computes the alignment between the atomic factual claims and various aspects expected to be presented in the output. We study three implementations of the ICAT framework, each with a different assumption on the availability of aspects and alignment method. By adopting data from the diversification task in the TREC Web Track and the ClueWeb corpus, we evaluate the ICAT framework. We demonstrate strong correlation with human judgments and provide comprehensive evaluation across multiple state-of-the-art LLMs. Our framework further offers interpretable and fine-grained analysis of diversity and coverage. Its modular design allows for easy adaptation to different domains and datasets, making it a valuable tool for evaluating the qualitative aspects of long-form responses produced by LLMs.
Beyond Factual Accuracy: Evaluating Coverage of Diverse Factual Information in Long-form Text Generation
Chris Samarinas1thanks: Equal contribution.    Alexander Krubner211footnotemark: 1thanks: Work done while visiting CIIR at UMass Amherst.    Alireza Salemi1 Youngwoo Kim1    Hamed Zamani1 1University of Massachusetts Amherst  2Salzburg University of Applied Sciences 1{csamarinas, asalemi, youngwookim, zamani}@cs.umass.edu 2akrubner.bin-m2022@fh-salzburg.ac.at
1 Introduction
The rapid advancement of Large Language Models (LLMs) has revolutionized long-form text generation, enabling increasingly sophisticated applications from report writing to complex question answering. However, this progress has highlighted a critical challenge: how do we effectively evaluate not just the factual accuracy of generated content, but also its completeness and coverage of diverse perspectives? While recent work has made significant progress in developing comprehensive evaluation frameworks for LLMs Li et al. (2024), existing metrics often focus on isolated aspects like factuality or response quality Min et al. (2023), failing to capture the multi-dimensional nature of high-quality long-form text.
The evaluation of long-form text generation presents unique challenges that go beyond traditional metrics Samarinas et al. (2024). Lexical overlap metrics like BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005) are fundamentally limited by their reliance on surface-level text similarity, making them inadequate for evaluating semantically equivalent but lexically different expressions. While more recent approaches like BERTScore Zhang et al. (2019) and G-Eval Liu et al. (2023) attempt to address this through semantic similarity, they still face fundamental limitations when applied to long-form content. These metrics cannot effectively verify factual accuracy or assess whether the response comprehensively covers all relevant aspects of a topic. The vast space of possible acceptable outputs makes it impractical to create comprehensive reference texts, leading to the development of reference-free evaluation methodologies like Prism Agrawal et al. (2021); Thompson and Post (2020). However, these approaches often struggle to effectively identify hallucinations and biases.
Recent approaches like FActScore Min et al. (2023) and VERISCORE Song et al. (2024) have addressed the factuality challenge by evaluating atomic claims against reliable sources, but factual accuracy alone is insufficient. A unified metric that considers both factuality and coverage is crucial not only for evaluation but also for optimizing LLM performance. Such a metric could serve as a reward function in reinforcement learning to simultaneously improve both the factual accuracy and comprehensive coverage of LLM outputs.
Consider a user asking “What are the health effects of coffee consumption?” While an LLM response might present entirely factual claims about coffee’s benefits, such as its role in improving alertness and potential protective effects against certain diseases, failing to address known health risks (like anxiety or sleep disruption) would present an incomplete and potentially misleading picture. This illustrates a critical gap in current evaluation approaches: the need to assess not just the accuracy of individual claims, but also the comprehensive coverage of diverse relevant information. This is particularly crucial for applications like medical information systems, policy analysis, or educational content where balanced, complete information is essential for informed decision making.
This paper introduces ICAT, a novel evaluation framework that addresses this gap by measuring both factual accuracy and coverage of diverse factual information in long-form text generation. Our key contributions include: 1) a modular evaluation framework that decomposes long-form text into atomic claims and evaluates both their factual accuracy and their alignment with expected aspects, 2) three implementations with varying degrees of automation, suitable for different evaluation scenarios, and 3) a comprehensive evaluation of various LLMs and demonstrating strong correlation with human judgments in a user study.
ICAT first breaks down the generated long text into atomic claims. Through retrieval from a (reliable) corpus $C$ or the Web, ICAT verifies each atomic claim to ensure its factuality. To measure completeness and coverage of diverse facts, ICAT requires a set of diverse aspects to compute an alignment between each atomic factual claim in the LLM response and the set of diverse aspects. We study three implementations of the ICAT framework with different assumptions on the availability of aspects and alignment methods.
In our experiments, we use ClueWeb The Lemur Project (2009) as the retrieval corpus. We solely focus on the English documents of the ClueWeb collection. In addition to custom open-source retrieval from ClueWeb, we explore web-based grounding using the Brave Search API. For experiments, we rely on the input queries from the TREC Web Track Clarke et al. (2009, 2010, 2011, 2012). The argument for this decision is based on the fact that TREC Web Track queries have also been used for search result diversification. This means that the queries include up to 7 aspects and documents are provided with aspect-level relevance annotations. Our experiments show that there is relatively strong corelation of ICAT with human judgments, showcasing the utility of this framework for evaluating coverage of diverse factual information in LLM responses without human input.
By offering a modular and adaptable framework, ICAT enables researchers to tailor the evaluation process to specific needs, making it a valuable tool for assessing the qualitative aspects of long-form responses produced by LLMs. The decomposition of LLM outputs into atomic claims and their alignment with specific topics makes the evaluation process highly interpretable - evaluators can trace exactly which claims support which topics and identify gaps in coverage. This granular analysis capability, combined with the framework’s ability to evaluate both factual accuracy and topic coverage, provides a more comprehensive assessment compared to existing metrics that only measure one of these aspects.
Figure 1: Retrieval-based evaluation of LLM responses with ICAT. Topic generation and coverage models are optional depending on the chosen evaluation method.
2 Related Work
Text Generation Evaluation Traditional approaches to evaluating generated text have primarily focused on n-gram overlap metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), and METEOR Banerjee and Lavie (2005). While these metrics are effective for assessing local coherence and fluency, they fail to capture higher-level aspects such as topic coverage and diversity. Recent work has introduced more sophisticated metrics like BERTScore Zhang et al. (2019), BLEURT Sellam et al. (2020), and unified multi-dimensional evaluators Zhong et al. (2022) which leverage pre-trained LLMs for more nuanced evaluation.
Topic Coverage and Diversity Research on evaluating topic coverage has roots in information retrieval, where metrics like $\alpha$-nDCG Clarke et al. (2008) and S-Recall were used to assess the topical diversity of search results. The concept of diversity in evaluating generated text can encompass various interpretations, including lexical diversity (analyzing the variety of words used) and topical diversity (assessing the range of topics covered). In the context of text generation, recent work has explored various approaches to measuring lexical diversity, including term overlap self-similarity such as Self-BLEU Zhu et al. (2018) and the proportion of distinct unigrams and bigrams in generated responses Li et al. (2016). However, research on evaluation of topical diversity in LLMs is currently limited.
LLM Evaluation Frameworks Several frameworks have been proposed for evaluating different aspects of LLM performance, including factuality Min et al. (2023); Song et al. (2024) and dialogue quality Mehri and Eskenazi (2020). Concurrent to this research, the AutoNuggetizer framework Pradeep et al. (2024) used LLMs to generate and assess the coverage of nuggets in text. The EXAM++ framework Farzi and Dietz (2024) evaluates information coverage by checking if system responses can answer query-related exam questions, focusing on high-level answerability. In contrast, our work provides more fine-grained explainability by aligning individual atomic claims with retrieved evidence to verify both factual accuracy and aspect coverage at a more granular level. This allows us to not only assess if an aspect is covered, but also identify the specific claims and evidence supporting that coverage.
Our work builds upon these foundations, specifically addressing the challenge of evaluating topic coverage in long-form text generation while considering factuality at the same time.
3 ICAT
Approach for Obtaining Diverse Aspects Approach for Claim-Aspect Alignment
ICAT-M Manual: ground-truth aspects Manual: retrieval-based method with aspect-level ground-truth alignment
ICAT-S Manual: ground-truth aspects Automatic: retrieval-based method with aspect-level LLM-based alignment
ICAT-A Automatic: LLM-based aspect generation Automatic: retrieval-based method with aspect-level LLM-based alignment
Table 1: The methods used for obtaining query aspects and claim-aspect alignment in each variant of ICAT.
Queries that require a long-form response, e.g., complex non-factoid questions, are often associated with multiple aspects. The response to these queries often include multiple claims, some of which may be factually accurate, while others may be inaccurate. An ideal response to these queries should not only contain factually accurate claims, but should also leave no aspect or perspective unaddressed. For instance, an ideal answer to a question about a legislation should cover perspectives from all political parties. An ideal answer to a question about the impact of a food or a medication on health should cover both positive, neutral, and negative perspectives. However, no existing evaluation metric can evaluate both factual accuracy and aspect coverage in long-form text generation. To address these, given a long output $y$ produced in response to an input $x$, the ICAT framework computes two main scores: factuality score and coverage score.
Factuality Score.
Building upon prior work, such as FActScore Min et al. (2023) and VERISCORE Song et al. (2024), Factuality Score measures the ratio (or percentage) of factually accurate claims in $y$. To do so, it is crucial that the generated claims are accurate. Let $AC(y)$ be a function that extracts atomic claims from a generated response $y$. Given the set of atomic claims $C=AC(y)$ made in $y$, we define the function $C_{T}=CG(C;K)$ that verifies the factuality of claims in $C$ using a given knowledge source $K$. Therefore, $C_{T}\subseteq C$ denotes the set of factually verified claims in response $y$. Factuality Score is then defined as follows:
$S_{\text{fact}}=\frac{|C_{T}|}{|C|}$
where ||$\cdot$|| ⋅ | denotes the cardinality of the given set.
Coverage Score.
To evaluate information coverage and diversity in $y$, Coverage Score measures the ratio of aspects being covered by the factually accurate claims in the given text. Hence, it is essential to identify which query aspects are accurately addressed in the generated response. Formally, coverage score can be defined as:
$S_{\text{coverage}}=\frac{|\{TO(c,K):c\in C_{T}\}\cap TQ(x)|}{|TQ(x)|}$
where $TO$ is a function that identifies the subtopics associated with claim $c$, and $TQ$ is a function that returns all aspects related to the input $x$. Note that aspect coverage is only computed for factually verified claims, i.e., $C_{T}$, instead of all claims. The reason is that non-factual claims should be avoided, regardless of the aspect they cover. Thus, they should not contribute to the coverage score.
The $\text{ICAT}_{\beta}$ Score.
Inspired by F-measure Van Rijsbergen (1979), we calculate the weighted harmonic average of these factuality and coverage scores, as follows:
$\text{ICAT}_{\beta}=(1+\beta^{2})\frac{S_{\text{fact}}\cdot S_{\text{coverage}}}{\beta^{2}S_{\text{fact}}+S_{\text{coverage}}}$
where parameter $\beta$ is a hyper-parameter that controls the trade-off between the factuality and coverage scores. In more detail, $\beta$ controls the weight of Coverage Score compared to Factuality Score. Thus, a higher $\beta$ signifies the impact of information coverage, while a lower $\beta$ prioritizes factual accuracy. The default value for $\beta$ is equal to $1$, where factuality and coverage score are weighted uniformly. Throughout this paper, when the value of $\beta$ is not explicitly mentioned, the default value of $1$ is being used.
Variants of ICAT
We study three variants of ICAT implementations based on how they obtain query aspects and compute the alignment between atomic claims and aspects. Table 1 describes the approaches used in these three variants and highlights their differences. ICAT-M assumes that a ground-truth set of diverse aspects are obtained manually and is available to the evaluation framework. It also assumes that ground-truth relevance annotations for each document in the corpus $C$ to each aspect are provided. Using this information, the retrieval model can identify which aspect is being covered by each atomic factual claim in the LLM response. ICAT-S similarly assumes that a ground-truth set of diverse aspects are obtained manually, however no aspect-level relevance judgment is available. Therefore, it uses an LLM to conduct pseudo-labeling and perform alignment between the atomic factual claims and the set of aspects. ICAT-A assumes that the aspect set is not available, so it first uses an LLM to automatically generate diverse aspects of the input and then conducts pseudo-labeling for alignment.
The rest of this section provides details on how to develop models for generating factual claims (i.e., function $AC$), how to validate the factuality of claims (i.e., function $CG$), and how to obtain query aspects and compute an alignment between factual claims and all query aspects (i.e., functions $TQ$ and $TO$). The rest of this section describes the approaches at high level to introduce the generic ICAT, while Section 4 provides the implementation details used in our experiments.
3.1 Atomic Claim Generation
The atomic claim generation process seeks to break down a given long text into standalone and atomic claim statements that preserve key context and maintain claim consistency (see Figure 2). The generated claims should strike an appropriate balance of granularity, ensuring they are self-contained and decontextualized. We assume the existence of a function $C=AC(y)$, which returns a set of atomic claims $C$, given the long output text $y$. There are various ways to implement this; one might consider each sentence or paragraph in $y$ as an atomic claim. However, this simple approach does not satisfy our expected self-containment and decontextualization qualities. Instead, we follow Min et al. (2023) and utilize an LLM $M_{\text{claims}}$ with the prompt shown in Figure 6. This prompt instructs the LLM to decompose the generated response $y$ into multiple self-explanatory and decontextualized sentences, each containing a single atomic fact. These sentences then constitute the set of atomic claims for the generated output, denoted as $C=AC(y)$. An example of this process is illustrated in Figure 2. An instruction-tuned LLM can be used as $M_{\text{claims}}$ in a zero- of few-shot setting; however, we found that smaller-scale LLMs (such as LLaMA Grattafiori et al. (2024) with 8 billion parameters) cannot accurately perform this task without fine-tuning. Therefore, we either use an LLM with higher capacity or distill knowledge into a smaller-scale LLM, enabling faster inference for our framework. The details of this distillation process are provided in section 4.1.
Figure 2: Example of atomic claim generation
3.2 Claim Grounding
To design the claim grounding function $CG$, for each claim $c\in C$, we employ a retrieval model $R$ to retrieve $n$ documents from the given knowledge source $K$. Subsequently, a natural language inference (NLI) model $M_{\text{NLI}}$ is used to determine whether the claim can be supported by any of the retrieved documents. If the claim can be inferred from at least one of the retrieved documents, it is considered grounded (i.e., validated, thus factually accurate); otherwise, it is not. The function returns a subset of $C$ that are found grounded.
3.3 Aspect Coverage Assessment
To calculate aspect coverage for some given input prompt (query) $x$, it is essential to have a list of diverse aspects for $x$ (i.e., $TQ(x)$) and a method to determine which aspect each claim pertains to (i.e., $TO(c,K):c\in C$).
Methods for Obtaining Diverse Query Aspects ($TQ$):
We propose two main methods to identify all aspects related to the query $x$:
  • Manual–Ground-truth Aspects: In this case, the aspects that should be included in the response to the query $x$ are provided as a reference for evaluation.
  • Automatic–LLM-based Aspect Generation: Building on previous work showing LLMs can effectively identify aspects of a query Samarinas et al. (2022), we use an LLM $M_{\text{subtopic}}$ with the prompt shown in Figure 7. This prompt instructs the LLM to generate up to 10 aspects for the query, covering the key aspects about it. This approach is useful when ground-truth aspects are unavailable.
Methods of Obtaining the Aspects of an Atomic Claim ($TO$):
We use two methods to identify the aspects related to an atomic claim $c$:
  • Manual–retrieval-based method with aspect-level ground-truth alignment: In this method, we assume access to a knowledge source $K$, where each document is annotated with the aspects it covers. To find the aspects that the claim $c$ covers, we use the retrieval model $R$ to retrieve $n$ documents. Then, according to the ranking, we find the first ranked document that supports the claim $c$ using the method in Section 3.2. The aspects of this document are considered as the aspects that the claim $c$ covers. If none of the documents support claim $c$, we assume that it does not cover any query aspect.
  • Automatic–retrieval-based method with aspect-level LLM-based alignment: In this method, we use an aspect-claim alignment LLM $M_{\text{coverage}}$ to determine which aspects each grounded claim covers. Given a query $x$, its aspects $TQ(x)$, and a set of grounded claims $C_{T}$, we prompt the LLM to analyze each claim and identify which aspects it addresses. The prompt (shown in Figure 5) instructs the LLM to output a structured mapping between claims and aspects, where each claim can be mapped to zero, one, or multiple aspects. This approach eliminates the need for aspect-level relevance judgments in the knowledge source while still maintaining a retrieval-based verification of factual accuracy. Unlike the manual method that assumes a claim covers the aspects associated with its supporting document, this method directly analyzes the semantic relationship between claims and aspects, leading to more accurate assessment.
Corpus-based Retrieval Web-based Retrieval
Method Coverage Model Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$ Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$
ICAT-M N/A 0.196 0.184 0.146 - - -
ICAT-S Llama-3.1-8B 0.247 0.255 0.200 0.294 0.252 0.204
Llama-3.1-70B 0.422 0.446 0.368 0.525 0.443 0.358
ICAT-A Llama-3.1-8B 0.161 0.125 0.097 0.033 0.021 0.014
Llama-3.1-70B 0.246 0.218 0.173 0.055 0.054 0.045
Table 2: Correlation of information coverage in the proposed evaluation methods with manual human annotations. Though ICAT-A shows weaker correlation than ICAT-S, we observed that auto-generated topics were more comprehensive and higher quality than the existing ones in the TREC dataset, suggesting ICAT-A would perform better with more complete ground-truth topics.
4 Implementation Details
4.1 Atomic Claim Generation
The claim generation module was trained with several key objectives in mind, building on recent work in atomic claim extraction Min et al. (2023). The model learned to extract standalone factual statements from text while maintaining factual consistency and simplifying complex statements. Special attention was paid to preserving important context and qualifiers, and generating claims at an appropriate granularity level Song et al. (2024). For this task we used Llama 3.1 8B Grattafiori et al. (2024) fine-tuned using QLoRA Dettmers et al. (2023) on synthetic examples. We tried using models of this size without fine-tuning, however we found that the generated claims are often not de-contextualized properly. Larger models with 70B or more parameters seem to be effective for this task without fine-tuning, however they are very expensive to run, especially for long texts.
The synthetic training data was generated through a multi-stage process by prompting Llama 3.1 405B. We began by generating 200 diverse high-level topics across multiple domains. For each topic, we generated 5 relevant entities. We then created variable-length paragraphs for each entity and generated the associated list of atomic claims for each of them. Using these 1000 synthetic examples, we fine-tuned the model for 1 epoch with batch size 16, learning rate 2e-4 and LoRA parameters $\alpha$ = 16 and rank = 64.
Our human evaluation comparing the fine-tuned 8B model with zero-shot 70B model revealed an interesting trade-off: while the 70B model achieves better semantic similarity with ground-truth facts, the fine-tuned 8B model produces better-formulated claims with higher fact precision (0.838 vs 0.753). This superior decontextualization is crucial for accurate retrieval-based grounding, as claims must be self-contained to match against corpus snippets effectively. This finding, combined with the significant computational cost savings (approximately 8.75× less memory and proportionally faster inference), justifies our design choice of using a fine-tuned smaller model. We plan to release a similarly fine-tuned adapter for fact-topic alignment to close the performance gap with the 70B model in that component as well. A detailed evaluation of the atomic claim generation quality is provided in Appendix A.5.
4.2 Topic Generation
Previous works have shown that LLMs can be very effective in query subtopic generation Samarinas et al. (2022). In our framework, for generating ground-truth topics given a query, we use the same base LLM as the one used in claim generation. Here we found that even without fine-tuning, Llama 3.1 8B can produce relevant topics. In order to reduce the need for extra resources to use a base and fine-tuned version of the LLM for claim generation, we use the VLLM library Kwon et al. (2023) to load the base model only once in memory and efficiently serve the adapter for the fine-tuned version. The quality evaluation of topic generation is detailed in Appendix A.5.
4.3 Claim Grounding
We implemented a two-stage approach for grounding atomic claims in the given text with a corpus. We first preprocess the corpus and generate chunks for each document with up to 128 words with 32 words overlap. We use a dense embedding model222hf.co/Snowflake/snowflake-arctic-embed-m Merrick et al. (2024) to produce embeddings for all snippets and FAISS Johnson et al. (2019) to build an efficient approximate nearest-neighbor index. We used IVF with HNSW for cluster assignment as our index type for fast search even when providing a large-scale corpus.
In the first stage, a retriever is used to obtain the $k=10$ most relevant snippets in the corpus for each claim. When web search is used instead of a corpus, we use the returned snippets from Brave Search API. In the second stage, a natural language inference model is used to filter only the supported claims. We use a model based on DeBERTa V3 He et al. (2021) fine-tuned on MultiNLI, FEVER and Adversarial NLI Williams et al. (2018); Thorne et al. (2018); Nie et al. (2020). A claim is kept if there is at least one snippet that supports it. Instead of using a LLM for filtering the supported claims, we use a much smaller BERT based model Devlin et al. (2019) fine-tuned specifically for this task.
We limit the snippet length because both NLI and dense embeddings models based on small pre-trained transformer LMs like BERT tend to have lower performance as the input length increases. The choice of using a specialized NLI model rather than a zero-shot LLM for claim verification was driven by both efficiency and effectiveness considerations. Our DeBERTa-based model achieves strong performance across multiple NLI benchmarks (MNLI: 0.903, FEVER-NLI: 0.777, ANLI-all: 0.579) while being orders of magnitude faster and more resource-efficient than LLM-based approaches. Given that grounding requires processing numerous claim-snippet pairs, efficiency is crucial for practical deployment and potential use as a reward function in reinforcement learning optimization of LLMs. A comprehensive evaluation of the NLI model performance across benchmarks is provided in Appendix A.5.
4.4 Aspect-Claim Alignment
Topic coverage is assessed using the same base LLM with claim and topic generation. Given a query, a list of enumerated atomic claims and a list of ground truth topics, the LLM is prompted to produce a list of covered topic ids with their associated claim ids in structured jsonl format. The accuracy of the aspect-claim alignment module is evaluated in Appendix A.5.
ICAT-M ICAT-S ICAT-A LLM $S_{\text{fact}}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-M}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-S}_{1}$ $S_{\text{coverage}}$ $\text{ICAT}\text{-A}_{1}$ Corpus-based Retrieval GPT-4 0.343 0.416 0.327 0.453 0.346 0.463 0.354 Llama-3-70B 0.327 0.451 0.335 0.464 0.355 0.466 0.354 Mixtral-8x22B 0.344 0.370 0.297 0.414 0.342 0.409 0.339 Openchat 3.5 (7B) 0.340 0.413 0.329 0.429 0.348 0.424 0.347 Web-based Retrieval GPT-4 0.748 - - 0.551 0.634 0.541 0.627 Llama-3-70B 0.714 - - 0.556 0.625 0.486 0.578 Mixtral-8x22B 0.749 - - 0.503 0.601 0.429 0.545 Openchat 3.5 (7B) 0.741 - - 0.520 0.611 0.444 0.555
Table 3: Evaluation of various LLMs using ICAT. Llama-3.1-70B is used for claim-aspect alignment.
5 Experimental Setup
Dataset.
We conducted our experiments using the ClueWeb09 Category B corpus–a large-scale web collection with over 50 million English documents The Lemur Project (2009). This corpus has been used in TREC Web Track from 2009 to 2012 Clarke et al. (2009, 2010, 2011, 2012), consisting of 200 topics, derived from a commercial search engine’s query log, balanced for popularity. Each topic includes a topic title (i.e., often used as the keyword search query), a description (i.e., detailed description of the information need), type, and subtopics (i.e., diverse aspects of the topic). The relevance judgments encompass 38,637 query-document pairs, with 19.06% (7366) marked as relevant. The dataset’s unique advantage lies in its comprehensive coverage of internet content and human-annotated relevance judgments for topical diversity assessment. Relevance was judged either binary or on a five-point scale (later converted to binary), with documents considered relevant when containing useful information for specific subtopics. In our experiments with this collection, we filtered out spam documents using the Waterloo spam scorer Cormack et al. (2011) with the threshold of 70%. We used BM25 to retrieve 1000 documents for each topic (given its title as the query string) and considered these documents for retrieval in our factual verification process. From the original 200 TREC topics, we filtered out approximately 21 queries that were not suitable for long-form text generation evaluation (e.g., queries asking for images, URLs, or specific data points), resulting in 179 valid queries. For human evaluation and detailed component analysis, we randomly sampled 50 queries from this filtered set due to cost constraints. These 50 queries were also used for LLM evaluation to ensure consistency and enable direct comparison with human judgments.
Experimental Setup.
We evaluated four state-of-the-art LLMs: GPT-4, Llama-3-70B-Instruct, Mixtral-8x22B-Instruct-v0.1, and Openchat 3.5 (a fine-tune of Mistral-7B) Wang et al. (2023). For each model, we generated responses for each test query. For the baselines in this paper, we used the query descriptions in their original format from the ClueWeb09 dataset as prompts, which are not optimized for producing very diverse outputs.
6 Experimental Results
Human Evaluation Study.
To validate ICAT’s effectiveness, we conducted a comprehensive human evaluation study using Amazon Mechanical Turk (AMT). For each query-answer pair, three independent annotators assessed the coverage of aspects through a custom interface (Figure 4 in Appendix A.4). We limited the HITs to adult workers from the US, UK, Australia and Ireland, with over 98% approval rate who have completed at least 5,000 assignments. The annotators were tasked with identifying whether specific aspects are present in a given LLM-generated text and highlighting corresponding text evidence for each identified aspect. To ensure quality annotations, we provided detailed guidelines with two reference examples. We use majority voting across annotators. The study achieved substantial inter-annotator agreement with Fleiss’s $\kappa$ = 0.829, which is considered as a substantial agreement. For each query, we calculated Coverage Scores based on the set of covered topics identified by each evaluation method (ICAT variants) and by human annotators, relative to the set of ground truth topics. These per-query coverage scores were then used to compute linear and rank-bsaed correlation metrics (i.e., Pearson’s $\rho$, Spearman’s $\rho$, and Kendall’s $\tau$) between the automated ICAT methods and human judgments.
The correlation analysis between ICAT variants and human judgments (see Table 2) reveals strong performance across most evaluation methods. Using Llama-3.1-70B as the coverage model, ICAT-S achieves the strongest correlations (Pearson’s $\rho$ = 0.422, p < 0.01; Spearman’s $\rho$ = 0.446, p < 0.01). Interestingly, ICAT-A demonstrates better performance than ICAT-M despite not requiring ground-truth aspects. While ICAT-A seems to have much weaker correlation than ICAT-S, it is worth noting that in practice, we observed the automatically generated topics to be higher quality and more exhaustive. Using a more comprehensive set of ground-truth topics, ICAT-A would probably demonstrate higher correlation with human judgements.
Comparison with Traditional Metrics for Text Generation.
To demonstrate the advantages of ICAT over traditional text generation evaluation metrics for assessing topic coverage, we compared our approach with widely-used metrics including BLEU, ROUGE, METEOR, and BERTScore. For this comparison, we concatenated the text of all ground-truth topics as a reference and evaluated the LLM responses against this reference. Table 4 presents the correlation results with human-annotated coverage scores.
Among traditional metrics, only BERTScore-recall shows statistically significant correlation with ground-truth coverage scores (Pearson’s $\rho$ = 0.291, p < 0.001). However, this correlation is still much lower than ICAT-S. Notably, n-gram based metrics (BLEU, ROUGE variants) show negative or near-zero correlations, demonstrating their inadequacy for evaluating topic coverage in long-form text. These results confirm that traditional metrics, designed primarily for surface-level text similarity, fail to capture the semantic coverage of diverse aspects that ICAT effectively measures.
Metric Pearson $\rho$ p-value Spearman $\rho$ p-value Kendall $\tau$ p-value
BERTScore-precision -0.003 0.966 0.048 0.499 0.019 0.709
BERTScore-recall 0.291 0.000 0.313 0.000 0.226 0.000
BERTScore-F1 0.120 0.090 0.188 0.008 0.126 0.013
METEOR 0.027 0.704 0.100 0.160 0.071 0.164
ROUGE-1 -0.058 0.415 0.016 0.819 0.014 0.785
ROUGE-2 -0.090 0.204 -0.032 0.655 -0.018 0.725
ROUGE-L -0.090 0.206 0.020 0.781 0.012 0.808
ROUGE-Lsum -0.090 0.206 0.020 0.781 0.012 0.808
BLEU -0.099 0.165 -0.059 0.409 -0.046 0.413
Table 4: Correlation of traditional text generation metrics with human-annotated topic coverage scores. Only BERTScore-recall shows statistically significant correlation, though still lower than ICAT variants.
Comparing factuality and coverage of information in state-of-the-art LLMs using ICAT.
Our experimental results reveal distinct patterns in how different LLMs balance factuality and coverage (see Table 3). When using corpus-based retrieval, Llama-3-70B demonstrates superior Coverage Score (0.451), while GPT-4 and Mixtral-8x22B show comparable factuality scores (0.343 and 0.344, respectively). However, Mixtral exhibits notably lower Coverage Score (0.370) compared to GPT-4’s 0.416, resulting in lower overall $\text{ICAT}_{1}$ scores (0.297 vs 0.327). Notably, when employing web-based retrieval, we observe substantially higher factuality scores across all models. This significant improvement can be attributed to the broader knowledge base available through web search, allowing more claims to be successfully grounded. Coverage scores also show improvement with web-based retrieval, though the increase is more modest. This suggests that while web search enables better fact verification, the comprehensiveness of aspect coverage is more dependent on the model’s capabilities than the retrieval source.
Controlling the trade-off between factuality and coverage using $\beta$.
Figure 3 illustrates how different values of $\beta$ affect the trade-off between factuality and coverage scores. Users can adjust $\beta$ based on their evaluation priorities: values of $\beta<1$ give more weight to factuality score, while $\beta>1$ emphasizes on coverage. In our experiments, for lower $\beta$ values, GPT-4 and Mixtral demonstrate superior performance in terms of $\text{ICAT}_{\beta}$. However, as $\beta$ increases and coverage becomes more important, GPT-4 and Llama achieve higher scores due to their stronger Coverage Score.
Component-level Evaluation.
To understand the contribution of each component in our framework, we conducted detailed evaluations of the atomic claim generation, topic generation, and coverage assessment modules. A comprehensive analysis of each component’s performance is provided in Appendix A.5.
Figure 3: ICAT-Aβ for various LLMs using Llama-3.1-70B as coverage model and the corpus as knowledge source. For low $\beta$ values where factuality has higher weight, GPT-4 and Mixtral have better performance while for higher values of $\beta$, GPT-4 and Llama have higher scores due to higher $S_{\text{coverage}}$.
7 Conclusions and Future Work
We presented ICAT, a comprehensive framework for evaluating topic coverage in LLM-generated text. Through extensive experimentation using the ClueWeb09 dataset, we demonstrated the framework’s effectiveness across different evaluation scenarios, with our best method achieving strong correlation with human judgments. The modular architecture of ICAT enables flexible adaptation to various evaluation requirements, from manual to automatic approaches for aspect identifications and alignment. Our results highlighted several key findings: (1) the importance of sophisticated coverage models in improving evaluation accuracy, (2) the viability of automatic evaluation approaches that maintain comparable performance to methods requiring ground truth annotations, and (3) the framework’s ability to provide meaningful assessments across different LLM architectures and scales. In future work, individual components of our ICAT could be improved in terms of effectiveness and efficiency. Additionally, exploring the relationship between model size, evaluation accuracy, and computational efficiency could provide valuable insights for practical applications. Last but not least, the potential bias introduced by using the same or similar LLM when generating ground-truth aspects should be investigated. By using our metric, other works can explore methods for optimizing LLMs to produce more comprehensive outputs.
Acknowledgments
This work was supported in part by the Center for Intelligent Information Retrieval (CIIR), in part by the Office of Naval Research contract number N000142212688, and in part by NSF grants #2143434 and #2106282. We acknowledge the support from the Austrian Marshall Plan Foundations, Stefan Wegenkittl, and Martin Uray who made Alexander Krubner’s visit to the CIIR possible. Any opinions, findings and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect those of the sponsors.
Limitations
Our evaluation framework, while showing promising results, suffers from several limitations that should be considered. First, our experiments reveal that even large language models with 70B parameters sometimes struggle with accurate aspect-claim alignment. This suggests that the correlation with human judgments could potentially be improved by specifically optimizing LLMs for this task, either through fine-tuning or more sophisticated prompting strategies.
Second, our current implementation uses zero-shot prompting for query aspect generation without systematic evaluation of this component’s effectiveness. Future work should explore methods to optimize and rigorously evaluate the aspect generation process, potentially through human evaluation or comparison with expert-curated aspect sets. This could lead to more reliable and comprehensive aspect coverage assessment.
In addition, there is a potential source of bias when using the same or similar LLM architecture both for generating query aspects and for producing responses for evaluation. This circular dependency might lead to artificially inflated performance metrics if the evaluated model shares similar biases or knowledge patterns with the model used for aspect generation. Future research should investigate the extent of this potential bias and explore methods to mitigate it, such as using diverse model architectures or more comprehensive human-curated aspects for evaluation.
Last but not least, our current evaluation is limited to English-language content using web-based corpora. This narrow focus excludes evaluation of multilingual capabilities and limits the framework’s applicability to other languages and cultures. Additionally, the reliance on web corpora may not be suitable for domains requiring specialized knowledge bases or authoritative sources. Future work should explore extending ICAT to support multilingual evaluation and integration with domain-specific knowledge bases to broaden its applicability across different languages, cultures, and specialized fields.
Appendix A Appendix
A.1 Atomic Claim Generation
Our atomic claim generation process transforms complex text into atomic, verifiable statements while preserving essential context, as shown in Figure 2. The process focuses on creating decontextualized, self-contained claims that each express a single verifiable fact. When breaking down complex sentences, the process maintains important qualifiers, conditions, and temporal information while replacing contextual references with their explicit referents. For instance, a complex sentence about coffee’s effects would be decomposed into separate claims about its components and their individual effects, with each claim being fully self-contained and independently verifiable.
A.2 Model Performance Analysis
The performance analysis presented in Tables 5 and 10 reveals several important patterns in model behavior. When comparing the 8B and 70B versions of Llama-3.1 as coverage models, we observe that the 8B model generally produces higher raw coverage scores, while the 70B model demonstrates stronger correlation with human judgments. This suggests a trade-off between computational efficiency and evaluation accuracy. The retrieval model comparison shows consistent advantages for dense retrieval methods over traditional BM25, with the Snowflake-Arctic-Embed models showing particular strength in handling queries where simple lexical matching is insufficient. Web-based retrieval consistently produces higher factuality scores compared to corpus-based approaches across all tested models. Table 10 analyzes the impact of different retrieval models on the correlation between ICAT variants and human judgments. The results show that larger dense retrieval models (Snowflake-Arctic-Embed-L) consistently outperform traditional BM25 across all ICAT variants, with improvements particularly notable in ICAT 2 and 3.
LLM Retrieval $S_{\text{fact}}$ ICAT-S ICAT-A
[…]
GPT-4 Corpus 0.343 0.563 0.394 0.563 0.394
Web Search 0.756 0.583 0.658 0.583 0.658
Llama-3-70B Corpus 0.327 0.616 0.401 0.616 0.401
Web Search 0.730 0.636 0.679 0.636 0.679
Mixtral-8x22B Corpus 0.344 0.567 0.401 0.567 0.401
Web Search 0.764 0.587 0.664 0.587 0.664
Openchat 3.5 (7B) Corpus 0.340 0.550 0.391 0.550 0.391
Web Search 0.752 0.570 0.648 0.570 0.648
Table 5: Comprehensive evaluation results comparing ICAT-S and ICAT-A variants using Llama-3.1-8B as the coverage model. Results show performance across different LLMs using both corpus-based and web-based retrieval methods, demonstrating consistent improvements in factuality and coverage scores when using web search compared to corpus-based retrieval. The table highlights how different retrieval methods affect both factuality ($S_{\text{fact}}$) and coverage scores across various model architectures.
Figure 4: Human annotation interface showing the supporting evidence highlights from 3 annotators for the given query in the shown LLM response.
A.3 Prompting Details
The framework employs three specialized prompts, each carefully designed for its specific task. The subtopic coverage assessment prompt (Figure 5) requests the identification of covered subtopics from a given list, requiring evidence in the form of fact numbers that explicitly appear in the text. The prompt specifies a structured JSON output format where each line contains a topic ID and its supporting evidence. The atomic claim generation prompt (Figure 6) focuses on extracting decontextualized, self-explanatory fact sentences from the input text, emphasizing the importance of resolved pronouns and independent context. The topic generation prompt (Figure 7) elicits possible subtopics or related queries for a given query, requiring them to be ordered by importance and formatted as JSON objects, with a maximum of 10 topics.
A.4 Human Evaluation Interface
The human evaluation interface shown in Figure 4 was developed to facilitate consistent and accurate topic coverage assessment. The interface presents the query and response in a split-screen layout, enabling annotators to highlight evidence for different topics using a color-coded system. To ensure annotation quality, we implemented minimum time requirements per response and included attention check questions. Annotators were guided to read responses completely before beginning their annotations and to identify text spans that provide direct evidence for topic coverage.
A.5 Component-level Evaluation
Atomic Claim Generation Quality. We evaluated the quality of atomic claim generation by comparing our fine-tuned Llama-3.1-8B model against the zero-shot Llama-3.1-70B model on 50 randomly sampled queries. Human annotators created ground-truth atomic facts for each query, and we assessed the generated claims using multiple metrics. To enforce local comparisons between atomic facts, we used a sliding window approach (window size ±3) when computing similarity metrics.
We evaluated using several metrics: (1) BERTScore333using microsoft/deberta-v3-base metrics for token-level similarity; (2) Semantic Precision/Recall/F1 computed using cosine similarities of sentence embeddings444using Alibaba-NLP/gte-base-en-v1.5, where precision measures how many generated claims have semantically similar ground-truth claims, recall measures coverage of ground-truth claims, and F1 is their harmonic mean; (3) Completeness, similar to semantic recall but with similarities binarized at threshold 0.8 to measure hard coverage; (4) traditional metrics (ROUGE, BLEU); and (5) Fact Precision, which measures the proportion of atomic claims that are properly formulated—specifically, claims that are decontextualized, self-contained, and verifiable in isolation without pronouns or missing context based on human judgements.
As shown in Table 6, the 70B model achieves higher scores on most semantic similarity metrics (e.g., BERTScore F1: 0.571 vs 0.499), suggesting better overall alignment with ground-truth facts. However, the fine-tuned 8B model demonstrates superior fact precision (0.838 vs 0.753), indicating better decontextualization and self-containment of claims—crucial properties for accurate retrieval-based grounding. This trade-off between semantic similarity and proper claim formulation justifies our choice of using a fine-tuned smaller model for efficient claim generation.
Metric Llama-3.1-8B Llama-3.1-70B Diff
BERTScore Precision 0.497 0.584 +0.088
BERTScore Recall 0.514 0.569 +0.055
BERTScore F1 0.499 0.571 +0.072
Semantic Precision 0.618 0.726 +0.109
Semantic Recall 0.808 0.825 +0.017
Semantic F1 0.681 0.761 +0.080
Completeness 0.537 0.593 +0.056
ROUGE-1 F 0.703 0.706 +0.003
ROUGE-2 F 0.555 0.563 +0.008
ROUGE-L F 0.599 0.610 +0.011
BLEU 0.379 0.391 +0.012
Fact Precision 0.838 0.753 -0.085
Table 6: Evaluation of atomic claim generation quality comparing fine-tuned 8B and zero-shot 70B models. The 8B model shows superior fact precision despite lower semantic similarity scores.
Topic Generation Quality. We assessed the relevance of automatically generated subtopics for the same 50 queries. Human annotators evaluated each generated subtopic as relevant or irrelevant to the query. Table 7 shows that both 8B and 70B models achieve remarkably high subtopic precision (0.966 and 0.958, respectively), demonstrating that even smaller models can effectively generate relevant query aspects when properly prompted. This validates our approach of using LLM-generated topics in ICAT-A when ground-truth aspects are unavailable.
Model Subtopic Precision
Llama-3.1-8B 0.966
Llama-3.1-70B 0.958
Table 7: Human evaluation of topic generation quality. Both models achieve high precision in generating relevant subtopics.
Coverage Assessment Accuracy. To evaluate the accuracy of our aspect-claim alignment module, we compared the coverage predictions of ICAT-S (using Llama-3.1-70B) against human judgments on the set of ground-truth topics. Table 8 shows strong agreement metrics, with high precision indicating that the topics identified as covered by our system are indeed covered according to human judgment, and good recall showing that most human-identified covered topics are captured by our system.
Coverage Metric Value
Coverage Precision 0.903
Coverage Recall 0.798
Coverage F1 0.835
Table 8: Accuracy of ICAT-S topic coverage predictions compared to human judgments using Llama-3.1-70B as the coverage model.
NLI Model Performance. Our claim grounding relies on a DeBERTa-based NLI model for efficiency. Table 9 shows the model’s performance across multiple NLI benchmarks. The strong performance on MNLI (0.903) and FEVER-NLI (0.777) demonstrates effectiveness on standard entailment and fact verification tasks. While performance on adversarial datasets (ANLI) is lower, as expected, the model still maintains reasonable accuracy even on the challenging ANLI-R3 subset (0.495). These results justify using this efficient model over more expensive LLM-based approaches for claim verification.
Dataset Accuracy
MNLI 0.903
FEVER-NLI 0.777
ANLI-all 0.579
ANLI-R3 0.495
Table 9: Performance of the DeBERTa-based NLI model used for claim grounding across multiple benchmarks.
Method Retrieval Model Pearson’s $\rho$ Spearman’s $\rho$ Kendall’s $\tau$
ICAT-M BM25 0.182 0.175 0.138
Snowflake-Arctic-Embed-M 0.196 0.184 0.146
Snowflake-Arctic-Embed-L 0.205 0.192 0.153
ICAT-S BM25 0.472 0.443 0.368
Snowflake-Arctic-Embed-M 0.489 0.452 0.376
Snowflake-Arctic-Embed-L 0.503 0.461 0.384
ICAT-A BM25 0.242 0.225 0.187
Snowflake-Arctic-Embed-M 0.251 0.236 0.191
Snowflake-Arctic-Embed-L 0.256 0.234 0.195
Table 10: Comparative analysis of retrieval model impact on ICAT variants’ correlation with human judgments, using Llama-3.1-70B for topic-claim alignment. Results demonstrate the superiority of dense retrieval models (Snowflake-Arctic-Embed) over traditional BM25 across all ICAT variants, with the largest improvements seen in ICAT-S and ICAT-A. The analysis includes three correlation metrics (Pearson’s, Spearman’s, and Kendall’s) to provide a comprehensive view of alignment with human assessments across different statistical measures.
Figure 5: Subtopic coverage prompt used to identify which subtopics are covered in a given text. The prompt instructs the model to analyze a list of atomic claims and return covered subtopics with supporting evidence in a structured JSON format. Each response must include topic IDs and corresponding fact numbers that explicitly appear in the text.
Figure 6: Claim generation prompt designed to extract atomic factual statements from input text. The prompt emphasizes the importance of creating decontextualized, self-explanatory sentences with resolved pronouns and independent context, ensuring each generated claim can stand alone as a verifiable statement.
Figure 7: Topic generation prompt used to automatically generate diverse subtopics for a given query. The prompt requests possible subtopics or related queries ordered by importance, with output formatted as JSON objects and limited to a maximum of 10 topics. This automated approach enables reference-free evaluation when ground-truth topics are unavailable.